OpenStudy (anonymous):

what is the x and y intercepts of 2x+3y=6

OpenStudy (amistre64):

divide it all by 6; then divide of the tops and read the intercepts from the bottoms

OpenStudy (amistre64):

\[2x+3y=6\] \[\frac{2x}{6}+\frac{3y}{6}=\frac{6}{6}\] \[\frac{x}{6/2}+\frac{y}{6/3}=1\] \[\frac{x}{3}+\frac{y}{2}=1\] when x=0, y=2 when x=3, y=0

OpenStudy (anonymous):

Visualize: The graph intercepts the X axis when Y=0, and intercepts the Y-axis when X=0. So, set x = 0, solve for Y. Set Y=0 solve for X. This trick will carry you through Calculus!

OpenStudy (anonymous):

An easy trick is to remember AX+BY=C. In this case, A=2, B=3 and C=6. Now, C/A is the x-intercept and C/B is the y-intercept.